OpenStudy (anonymous):

8b + 12 = 5(4 + 2b) 8b = 8 + 10b −2b = 8 b = −4 8b + 12 = 20 + 2b 8b + 12 = 20 + 10b 8b + 12 = 9 + 7b 8b + 12 = 4 + 10b

OpenStudy (blurbendy):

what's the question?

OpenStudy (anonymous):

: A partial solution to the equation 8b + 12 = 5(4 + 2b) is given below. Choose the step that is missing from the solution.

OpenStudy (blurbendy):

8b + 12 = 20 + 10b

OpenStudy (anonymous):

so the second one

OpenStudy (blurbendy):

yes

OpenStudy (blurbendy):

5(4 + 2b) = 20 + 10b